<p>09:34 - Moved to Run 9 in preparation of moving detectors. No target yet. Beam in the ring at 750 keV/A.</p>
<p>Temperatures OK (attach)</p>
<p>System wide checks OK except 16 fails clock and ADC calib</p>
<p>Rates OK (attach)</p>
<p>Writing to disk OK (attach)</p>
<p>Resetting the sorting script. Moving detectors.</p>
<p>Moved detectors in. See attached. Stripe / shadow quite clear. Some oddities on the top-right DSSD. May be due to ExEy cuts and multiplicity cuts.</p>
<p>Moved on tun run 9 to acquire some stats without motion.</p>
<p> </p>
<p>10:13 Temperatures still OK. Rates OK. Leakage currents stable.</p>
<p>We were cutting the beam.</p>
<p>Increased repetition to x5.</p>
<p>Motors fully out - run 10 started.</p> |